### v2

# Ultrasonic Medical Tissue Imaging Using Probabilistic Inversion: Leveraging Variational Inference for Speed Reconstruction and Uncertainty Quantification

Abstract
Full Waveform Inversion (FWI) is a promising technique for achieving high-resolution imaging in medical ultrasound. Traditional FWI methods suffer from issues related to computational efficiency, dependence on initial models, and the inability to quantify uncertainty. This study introduces the Stein Variational Gradient Descent (SVGD) algorithm into FWI, aiming to improve inversion performance and enhance uncertainty quantification. By deriving the posterior gradient, the study explores the integration of SVGD with FWI and demonstrates its ability to approximate complex priors. In-silico experiments with synthetic data and real-world breast tissue data highlight the advantages of the SVGD-based framework over conventional FWI. SVGD-based FWI improves inversion quality, provides more reliable uncertainty quantification, and offers a tighter bound for the prior distribution. These findings show that probabilistic inversion is a promising tool for addressing the limitations of traditional FWI methods in ultrasonic imaging of medical tissues.
1 Introduction
Ultrasound imaging plays an important role in diagnostics and monitoring in medical applications, providing real-time, non-invasive images of the internal body structures [1, 2, 3, 4]. Despite its widespread applications, traditional B-mode ultrasound imaging has limited spatial resolution, making it difficult to clearly display complex tissue structures and heterogeneous media [5, 6]. Moreover, the B-mode imaging is highly susceptible to noise and artifacts, further compromising its diagnostic accuracy and robustness.
In recent years, Full Waveform Inversion (FWI) has emerged as an advanced technique to enhance ultrasound imaging quality by utilizing the full ultrasonic waveform, rather than relying solely on amplitude or time-of-flight data [7, 8, 9]. FWI iteratively adjusts model parameters to minimize the discrepancy between observed and simulated data, achieving high-precision estimations of medium properties [10]. FWI has been proven to be both feasible and effective for inversing the speed of sound (SOS) image of musculoskeletal tissue [11, 12, 13] and brain tissues [14], addressing challenges such as image distortion and low signal-to-noise ratio. In quantitative imaging of bone tissues, FWI is capable of accurately reconstructing the fine structure of bone joints and capturing the cortical bone thickness [15].
Traditional FWI algorithms face several challenges, including low computational efficiency, strong dependence on initial models, and difficulties in accurately quantifying parameter uncertainty [16]. In uncertainty estimation for FWI, Markov Chain Monte Carlo (MCMC) methods, such as the Metropolis-Hastings algorithm [17], have traditionally been used to sample from complex posterior distributions [18, 19]. However, MCMC methods are inefficient in high-dimensional spaces due to their random-walk behavior, which leads to slow convergence and the tendency to become trapped at individual maxima in the probability distribution [20]. Although more advanced MCMC techniques, such as reversible-jump MCMC [21, 22, 23], stochastic Newton MCMC[24, 25], and Hamiltonian Monte Carlo [26, 27], aim to improve efficiency, these methods remain intractable for FWI due to their extremely high computational cost.
To improve the efficiency of MCMC methods, Variational Inference (VI) techniques are increasingly applied to estimate the posterior distribution of model parameters and address uncertainties inherent in the inversion process [28, 29, 30]. Instead of directly calculating the posterior distribution, VI approximates it by minimizing the Kullback-Leibler (KL) divergence between the true posterior and a simpler, tractable distribution [31, 32]. Recent research, such as in stochastic variational inference (SVI) for medical ultrasound imaging [33], has demonstrated how gradient-based methods can efficiently solve inverse problems. The derivation of SVI depends on selecting a suitable prior distribution, which can be challenging to determine in practice. Stein Variational Gradient Descent (SVGD), a VI-based method, uses a set of particles to approximate more complex prior distributions, such as a Gaussian mixture model [34]. The SVGD approach enhances both the efficiency and accuracy of VI in high-dimensional inverse problems, making it a promising method for addressing challenges in ultrasound imaging.
This paper explores the application of SVGD in medical ultrasound imaging, focusing on enhancing the efficiency and robustness of the inversion process. The core principles of the wave equation used in FWI to model wave propagation are outlined, followed by a discussion of the challenges in model inversion due to noise and complex data. The paper then introduces VI as a method for incorporating uncertainty into the inversion process, improving upon traditional optimization techniques. SVGD is presented as an advanced VI-based algorithm that optimizes the approximation of posterior distributions. The gradient of the posterior distribution is derived, and SVGD is applied to solve inverse problems in medical ultrasound imaging. Through simulations using both synthetic and real-world medical data, the results demonstrate that SVGD leads to faster convergence, improves image reconstruction quality, and provides more reliable uncertainty quantification.
This paper presents a theoretical framework for applying the SVGD algorithm to FWI. Section 2 introduces the foundational concepts of FWI and VI, providing the necessary background to understand their integration within a probabilistic framework. Section 3 focuses on the application of SVGD to FWI, including the derivation of the posterior gradient and an example demonstrating SVGD’s ability to approximate Gaussian Mixture Model. In Section 4, the framework is applied to synthetic data generated from linear-array and ring-array transducers. Section 5 extends the methodology to breast tissue data. Both Sections 4 and 5 use conventional FWI as a baseline to evaluate the inversion quality of SVGD. Section 5 also employs the Stochastic Variational Inference (SVI) for a comparative analysis of the uncertainty quantification provided by SVGD.
2 Probabilistic Inversion in FWI: Bayesian Framework and Variational Approximations
2.1 Acoustic Wave Propagation and Data Generation in FWI
The wave equation in Full Waveform Inversion (FWI) is used to simulate wave propagation and iteratively update the model to match the observed data. The most commonly used wave equation is the acoustic wave equation [35, 36],
$\frac{1}{\bm{c}^{2}}\frac{\partial^{2}\bm{u}(\bm{x},t)}{\partial t^{2}}-\nabla^{2}\bm{u}(\bm{x},t)=\bm{s}(\bm{x},t),$ (1)
where $u(\bm{x},t)$ represents the wavefield, such as pressure or displacement, at position $\bm{x}=(x_{1},x_{2},\cdots,x_{d})$ and time $t$. The notation $\bm{c}=\bm{c}(\bm{x})$ is the speed of sound (SOS), which varies spatially in a heterogeneous medium. The operator $\nabla^{2}=\sum_{i}\partial^{2}/\partial x_{i}^{2}$ is the Laplacian operator, representing the spatial derivatives of the wavefield. The source signal $\bm{s}(\bm{x},t)$ describes the spatial and temporal distribution of the energy input into the system. The exact form of $\bm{s}(\bm{x},t)$ depends on the specific modeling of the ultrasound source. For a simple point source, the expression is
$\bm{s}(\bm{x},t)=\delta(\bm{x}-\bm{x}_{s})\bm{f}(t),$
where $\delta(\bm{x}-\bm{x}_{s})$ is the Dirac delta function, representing a point source located at position $\bm{x}_{s}$. The function $\bm{f}(t)$ is a time-dependent function that models the temporal profile of the source, such as a pulse or continuous wave.
The receivers, positioned at $\bm{x}_{r}$, convert the mechanical vibrations of the incoming acoustic waves into electrical signals. These signals consist of both reflected and transmitted waves. To predict the wavefield data recorded by the receivers, the wave equation (1) is numerically solved using the finite difference method. The obtained data $\bm{d}_{r}$ is a time series and described by a functional relation that involves the position $\bm{x}_{r}$, time $t$, and material properties $\bm{m}$ of the medium. This relationship is
$\bm{d}_{r}=\mathcal{L}(\bm{x}_{r},t,\bm{m}),$ (2)
where $\mathcal{L}$ is a nonlinear operator that models the conversion of the mechanical displacement (or wavefield) into electrical signals by the transducer.
While the predicted data $\bm{d}_{r}$ is obtained by solving the wave equation with the parameters $\bm{m}$, real-world measurements $\bm{d}_{r}^{\text{obs}}$ typically exhibit discrepancies when compared to $\bm{d}_{r}$. The discrepancies are modeled as
$\bm{d}_{r}^{\text{obs}}=\mathcal{L}(\bm{x}_{r},t,\bm{m})+\bm{\epsilon},$ (3)
where $\bm{\epsilon}$ captures errors arising from an un-converged SOS model, as well as factors including sensor imperfections, environmental conditions, and inherent measurement noise. Therefore, the observed data $\bm{d}_{r}^{\text{obs}}$ is often a noisy and potentially distorted version of the predicted data $\bm{d}_{r}$, and addressing these discrepancies is essential for accurate model inversion and updating in FWI.
2.2 Misfit Function and Optimization in FWI
FWI is a powerful computational technique used to create high-resolution models of tissue properties. Unlike traditional B-mode methods that rely on simplified approximations, FWI makes use of the full information contained in the entire waveform of the received signal, including both amplitude and phase information. The primary objective of FWI is to iteratively update the model parameters $\bm{m}$ by comparing the experimentally observed data $\bm{d}_{r}^{\text{obs}}$ with the predicted data $\bm{d}_{r}$.
FWI begins with an initial model of the medium, which serves as an approximation of the model parameters $\bm{m}$ at each spatial point. Using this initial model, the predicted wavefield data $\bm{d}_{r}$ is computed numerically by solving the wave equation (1). The core of FWI is the iterative updating of the parameters $\bm{m}$ by minimizing the difference between $\bm{d}_{r}^{\text{obs}}$ and $\bm{d}_{r}$. The difference between the two datasets is quantified using a misfit function, often based on the least-squares error. The most commonly used misfit function is the $l_{2}$ norm, which measures the sum of squared differences across all receivers,
$l_{2}(\bm{d}_{r}^{\text{obs}},\bm{d}_{r})=\frac{1}{2}\sum_{r}\|\bm{d}_{r}-\bm{d}_{r}^{\text{obs}}\|^{2}.$ (4)
The goal of FWI is to find the optimal parameters $\bm{m}^{*}$ that minimize the misfit function, i.e., to solve the following optimization problem:
$\bm{m}^{*}=\mathop{\arg\min}_{\bm{m}}\ l_{2}(\bm{d}_{r}^{\text{obs}},\bm{d}_{r}).$ (5)
In practice, this minimization is achieved through iterative optimization techniques. The optimization process is typically nonlinear, as the relationship between the model parameters and the observed data is often complex and involves multiple reflections, scattering, and other high-order wave phenomena.
One of the key challenges of FWI is the possibility for the optimization to converge to local minima, especially when the initial model is far from the true solution. To mitigate this issue, careful initialization and regularization techniques are often employed. Additionally, FWI is computationally expensive, as both the forward and adjoint problems need to be solved repeatedly during the inversion process. Despite these challenges, FWI provides a powerful framework for obtaining high-resolution, accurate models of the medium’s properties, which significantly improve imaging and characterization in applications.
2.3 Variational Inference Solves Bayes’ Equation
In recent years, Variational Inference (VI) has emerged as a powerful tool for improving the efficiency and robustness of inversion techniques. VI provides an alternative approach to classical optimization by approximating complex posterior distributions with simpler, tractable distributions, thus reducing the computational burden. The goal of FWI is to find the optimal parameters $\bm{m}^{*}$ satisfying Eq. (5), VI treats the parameters $\bm{m}$ as random variables and infers their distribution given the observed data. This allows for more reliable reconstructions of tissue properties, providing both the best estimate of model parameters and a measure of uncertainty that is crucial for clinical decision-making. By approximating the true posterior distribution through a variational family, VI facilitates efficient model updating and allows for more robust handling of noisy or incomplete data. This is beneficial in medical ultrasound imaging, where measurement noise, sensor limitations, and complex tissue structures often make traditional deterministic inversion methods prone to local minima or overfitting. VI-based methods can enhance the accuracy and stability of FWI, providing more reliable reconstructions of tissue properties.
Bayes’ theorem provides the foundational framework for VI-based methods, which update the parameters $\bm{m}$ based on observed data $\bm{d}^{\text{obs}}$. Bayes’ theorem expresses the posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$ as
$p(\bm{m}|\bm{d}^{\text{obs}})=\frac{p(\bm{d}^{\text{obs}}|\bm{m})p(\bm{m})}{p(\bm{d}^{\text{obs}})},$ (6)
where $p(\bm{m})$ is the prior distribution that encodes our initial beliefs about the model parameters before observing the data. The evidence distribution $p(\bm{d}^{\text{obs}})$ (also known as marginal likelihood) acts as a normalizing constant ensuring that the posterior distribution integrates to one. The likelihood distribution $p(\bm{d}^{\text{obs}}|\bm{m})$ represents the probability of observing the data given a set of model parameters.
Computing the exact posterior distribution in Eq. (6) often requires solving complex integrals, which can be computationally intractable. To illustrate this, consider the joint distribution between the prior and the likelihood, i.e., $p(\bm{m},\bm{d}^{\text{obs}})=p(\bm{d}^{\text{obs}}|\bm{m})p(\bm{m})$. The evidence $p(\bm{d}^{\text{obs}})$ is the marginal of the joint distribution $p(\bm{m},\bm{d}^{\text{obs}})$,
[…]
The numerical integrate the integral in Eq. (7) is often challenging due to the high dimensionality of the model parameters $\bm{m}$. As a result, directly computing the posterior becomes intractable.
Another widely used numerical method to solve the posterior distribution in Eq. (6) is Markov Chain Monte Carlo (MCMC) [17, 37]. MCMC is a class of algorithms that generates samples from a probability distribution, allowing for an approximation of the posterior distribution. MCMC constructs a Markov chain, and the distribution of the chain’s states approaches the target posterior distribution after a large number of iterations. The Monte Carlo estimate of the posterior distribution $\hat{p}(\bm{m}|\bm{d}^{\text{obs}})$ is obtained by averaging over the samples generated by the chain,
$\hat{p}(\bm{m}|\bm{d}^{\text{obs}})=\frac{1}{c}\sum_{i}p(\bm{d}^{\text{obs}}|\bm{m}^{i})p(\bm{m}^{i}),$ (8)
where $p(\bm{m}^{i})$ represents the prior distribution evaluated at the $i$-th sample. The notation $c$ is the normalizing constant ensuring that the estimated posterior integrates to one. Although MCMC is a useful tool, it is computationally intensive, requiring numerous iterations to achieve accurate results. Besides, MCMC may struggle with high-dimensional spaces or complex dependencies between parameters.
The posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$ is often computationally challenging to obtain using analytical integral or numerical MCMC, which limits their applicability in FWI. To address these limitations, VI approximates the posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$ by minimizing the KL divergence between the true posterior and a simpler, tractable variational distribution $q(\bm{m})$. The variational distribution $q(\bm{m})$ belongs to a chosen family of distributions $\mathcal{Q}$, and the optimal distribution $q^{*}(\bm{m})$ is obtained by
$q^{*}(\bm{m})=\mathop{\arg\min}_{q(\bm{m})\in\mathcal{Q}}D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right],$ (9)
where the KL divergence is defined as
[…]
The KL divergence is non-negative and measures the difference between two probability distributions [38, 39].
[…]
$D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=E_{q}\left[\log(q(\bm{m}))\right]-E_{q}\left[\log(p(\bm{m}|\bm{d}^{\text{obs}}))\right].$ (11)
Substituting Bayes’ theorem (6) into Eq. (11),
$D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=E_{q}\left[\log(q(\bm{m}))\right]-E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]+E_{q}\left[\log(p(\bm{d}^{\text{obs}}))\right].$ (12)
In the above equation, the evidence $\log(p(\bm{d}^{\text{obs}}))$ is difficult to compute directly. Consequently, minimizing the KL divergence cannot be done directly. Rearranging the Eq. (12) gives
$E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]-E_{q}\left[\log(q(\bm{m}))\right]=\log(p(\bm{d}^{\text{obs}}))-D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right].$ (13)
Since $D_{\text{KL}}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]\geq 0$, a lower bound for the evidence is given by
$\log(p(\bm{d}^{\text{obs}}))\geq E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]-E_{q}\left[\log(q(\bm{m}))\right].$ (14)
The right-hand defines the Evidence Lower Bound (ELBO), denoted by
$\text{ELBO}(q)\triangleq E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]-E_{q}\left[\log(q(\bm{m}))\right].$ (15)
Substituting the definition of ELBO into Eq. (13) yields
$\text{ELBO}(q)=\log(p(\bm{d}^{\text{obs}}))-D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right].$ (16)
Because $\log(p(\bm{d}^{\text{obs}}))$ is constant for a specific problem, minimizing the KL divergence in Eq. (9) is equivalent to maximizing the ELBO. The optimization problem in Eq. (9) becomes
$q^{*}(\bm{m})=\mathop{\arg\max}_{q(\bm{m})\in\mathcal{Q}}\text{ELBO}(q).$ (17)
The definition of $\text{ELBO}(q)$ does not involve the intractable term $p(\bm{d}^{\text{obs}})$, maximizing the ELBO is computationally easier than directly minimizing the KL divergence. Observing Eq. (14) and (16),
$\log(p(\bm{d}^{\text{obs}}))=\text{ELBO}(q)\ \text{if and only if}\ D_{\text{KL}}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=0.$
The maximal value of $\text{ELBO}(q)$ is achieved when the approximating distribution $q(\bm{m})$ is exactly equal to the posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$.
By deriving the ELBO, VI reduces the computational burden compared to exact Bayesian inference methods, such as analytically solving Eq. (7) or using MCMC methods. The flexibility of VI lies in the choice of optimization techniques and the variational family $\mathcal{Q}$, which leads to the development of various algorithms. Among these, Stein Variational Gradient Descent (SVGD) stands out as a flexible and efficient method. In the following, this paper introduces SVGD and demonstrates its efficiency in solving inverse problems.
3 SVGD-Based Posterior Approximation and Uncertainty Quantification in FWI
3.1 Overview of SVGD
SVGD is an advanced algorithm designed for approximating complex posterior distributions in VI[34]. The SVGD algorithm forms a natural counterpart to gradient descent for optimization problems. Instead of relying on a fixed set of particles, SVGD actively transports particles to better match the target distribution by leveraging functional gradient descent.
The core idea of SVGD is to apply an incremental transformation to the model parameters $\bm{m}$. This transformation is defined by the identity map [34]
$f(\bm{m})=\bm{m}+\epsilon\,\bm{g}(\bm{m}),$ (18)
where $\bm{g}(\bm{m})$ is a smooth function that represents the perturbation direction, and $\epsilon$ is a scalar that controls the magnitude of the perturbation. This transformation is designed to be an invertible map, satisfying the inverse function theorem, and gradually updates the approximate distribution to match the true posterior.
The density of the transformed distribution is denoted as $q_{f}(z)$, where $z=f(\bm{m})$ is the transformed model. To optimize the approximation, SVGD minimizes the KL divergence between the transformed distribution $q_{f}(z)$ and the true posterior $p(\bm{m}|\bm{d}^{\text{obs}})$. The gradient of the KL divergence with respect to $\epsilon$ is
$\left.\nabla_{\epsilon}D_{KL}\left[q_{f}\|p\right]\right|_{\epsilon=0}=-E_{q}\left[\operatorname{trace}\left(\mathcal{A}_{p}\bm{g}(\bm{m})\right)\right],$ (19)
where $\mathcal{A}_{p}$ is the Stein operator, expressed as
$\mathcal{A}_{p}\bm{g}(\bm{m})=\nabla_{\mathrm{m}}\log p\left(\bm{m}\mid\bm{d}_{\mathrm{obs}}\right)\bm{g}(\bm{m})^{T}+\nabla_{\mathrm{m}}\bm{g}(\bm{m}).$ (20)
Eq. (19) describes how the perturbation function $\bm{g}(\bm{m})$ is updated to reduce the KL divergence between the approximate distribution and the true posterior. The gradient of the KL divergence informs the direction to move the model parameters in order to improve the approximation.
To find the optimal perturbation direction, Ref. [34] derives the steepest descent that maximizes the negative gradient in Eq. (19) as
$\bm{g}^{*}(\bm{m})=E_{\bm{m}\sim q}\left[k(\bm{m},\bm{m}^{\prime})\nabla_{\bm{m}}\log p\left(\bm{m}\mid\bm{d}_{\mathrm{obs}}\right)+\nabla_{\bm{m}}k(\bm{m},\bm{m}^{\prime})\right],$ (21)
where $k(\bm{m},\bm{m}^{\prime})$ is a kernel function that quantifies the similarity between model parameters $\bm{m}$ and $\bm{m}^{\prime}$. The kernel function determines how the perturbation at one point influences nearby points in the model space. One common choice for the kernel function is the Radial Basis Function (RBF) kernel,
$k(\bm{m},\bm{m}^{\prime})=\exp\left(-\frac{\|\bm{m}-\bm{m}^{\prime}\|^{2}}{2h^{2}}\right),$ (22)
where $h$ is a hyperparameter that controls the width of the kernel and determines the influence of one point on another. The RBF kernel is widely used due to its smoothness and locality, ensuring that points that are closer in model space will have more influence on each other. The gradient of the kernel function with respect to $\bm{m}$ is
$\nabla_{\bm{m}}k(\bm{m},\bm{m}^{\prime})=-\frac{\bm{m}-\bm{m}^{\prime}}{\sigma^{2}}\exp\left(-\frac{\|\bm{m}-\bm{m}^{\prime}\|^{2}}{2\sigma^{2}}\right).$ (23)
The final update rule at each iteration is
$\bm{m}_{i}^{t+1}=\bm{m}_{i}^{t}+\epsilon^{t}\left[k\left(\bm{m}_{i}^{t},\bm{m}_{j}^{t}\right)\nabla_{\bm{m}_{i}^{t}}\log p\left(\bm{m}_{i}^{t}\mid\bm{d}^{\mathrm{obs}}\right)+\nabla_{\bm{m}_{i}^{t}}k\left(\bm{m}_{i}^{t},\bm{m}_{j}^{t}\right)\right].$ (24)
The above iterative update process, using the kernel function and the gradient of the log posterior, refines the approximation to the true posterior distribution.
As demonstrated in Ref. [34], SVGD effectively employs particles to approximate the mean and standard deviation of complex distributions, including Gaussian Mixture Model (GMM). GMMs are widely used to represent complex probability distributions due to their flexibility in capturing multimodal behaviors. The GMM distribution $p(x)$ is defined as a weighted sum of Gaussian components,
$p(x)=\sum_{i}w_{i}p_{i}(x)=\sum_{i}w_{i}\mathcal{N}(x;\mu_{i},\sigma_{i}),$ (25)
[…]
$p_{i}(x)=\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(-\frac{(x-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right).$ (26)
The mean of the overall GMM is calculated as
$\mu_{\text{GMM}}=\sum_{i}w_{i}\mu_{i},$ (27)
which reflects the weighted contribution of each component’s mean. The standard deviation of the GMM is
$\sigma_{\text{GMM}}=\sqrt{\sum_{i}w_{i}\left(\sigma_{i}^{2}+\mu_{i}^{2}\right)-\mu_{\text{GMM}}^{2}}.$ (28)
The formulation (28) captures the combined spread of the distribution, accounting for both the variances of individual components and the deviation of their means.
Figure 1: SVGD approximates the mean and standard deviation of the Gaussian Mixture Model (GMM). (a) The black line represents the density distribution of the GMM, with mean $\mu_{\text{GMM}}=0.9$ and standard deviation $\sigma_{\text{GMM}}=1.446$. The blue, orange, and green lines correspond to the three components of the GMM, with weights $w_{1}=0.2$, $w_{2}=0.5$, and $w_{3}=0.3$, respectively. (b) The initial distribution of the particles, generated from the standard normal distribution $\mathcal{N}(x;0,1)$. (c) The results of the SVGD algorithm. The dashed red line shows the ground truth mean. The mean and standard deviation of the particles are $0.883$ and $1.314$, respectively. (d) The MSE of the mean and standard deviation plotted against the iterations of the SVGD algorithm.
Mathematically, any continuous probability distribution can be approximated arbitrarily well by a GMM given a sufficient number of components [40]. However, for complex target distributions, optimizing the means, variances, and weights of the Gaussian components becomes computationally expensive as the number of components increases. SVGD offers a nonparametric approach to approximate distributions by iteratively moving particles to better align with the target distribution. SVGD is advantageous for scenarios where the target distribution is intractable, as it eliminates the need for explicit parameter optimization. Consequently, if SVGD effectively approximate a GMM, it could also be leveraged to approximate other complex distributions with better efficiency. In the following, an example is employed to illustrate the effectiveness of SVGD in approximating the mean and standard deviation of a GMM.
Assuming a GMM with three components, the model is defined as
$p(x)=0.2\mathcal{N}(x;-1,1)+0.5\mathcal{N}(x;1,0.5^{2})+0.3\mathcal{N}(x;2,1.5^{2}),$ (29)
and the corresponding density and its components are illustrated in Figure 1 (a). By applying Eqs. (27) and (28), the mean and standard deviation of the GMM are
$\mu_{\text{GMM}}=0.9,\ \sigma_{\text{GMM}}=1.446,$ (30)
which represent the ground truth values for the SVGD algorithm to approximate. The SVGD algorithm begins with $n$ particles generated from the standard Gaussian distribution $\mathcal{N}(x;0,1)$, as depicted in Figure 1 (b). The initial random distribution of the particles shows noticeable differences from the GMM distribution shown in Figure 1 (a). To apply the SVGD algorithm to the GMM, the approximated mean and standard deviation are denoted as $\hat{u}$ and $\hat{\sigma}$, respectively, and the fit is evaluated using the mean squared error (MSE) between the true and approximated values. Following the SVGD update steps in Eq. (24), the final distribution of the particles is shown in Figure 1 (c), where most particles are centered around the ground truth mean. The computed mean and standard deviation of the particles are $\hat{u}=0.883$ and $\hat{\sigma}=1.314$, respectively, which are close to the ground truth values. Figure 1 (d) displays the MSE of the mean over iterations, with the inset showing the variations in the MSE of the standard deviation. After 600 iterations, both the MSE of the mean and standard deviation converge to their optimal values.
3.2 Gradient Calculation of Log-Posterior for SVGD Method
To apply the SVGD update rule in FWI, by substituting Bayes’ theorem (6) into the gradient of the log-posterior (24), yields
$\nabla_{\bm{m}}\log p\left(\bm{m}\mid\bm{d}^{\mathrm{obs}}\right)=$ $\nabla_{\bm{m}}\log p\left(\bm{d}^{\mathrm{obs}}\mid\bm{m}\right)+\nabla_{\bm{m}}\log p\left(\bm{m}\right).$ (31)
In the process of FWI, the likelihood term $p(\bm{d}^{\text{obs}}|\bm{m})$ is typically modeled using a misfit function, such as the least-squares error (4). Using this least-squares error formulation, the Gaussian likelihood function is expressed as [33],
$p\left(\bm{d}^{\mathrm{obs}}\mid\bm{m}\right)=\operatorname{trace}(2\pi\mathbf{I})^{-\frac{1}{2}}\exp\left\{-\frac{1}{2}\sum_{r}\|\bm{d}_{r}-\bm{d}_{r}^{\text{obs}}\|^{2}\right\}.$ (32)
where $\mathbf{I}$ is the identity matrix. The exponential term in Eq. (32) cancels with the logarithm,
$\log p\left(\bm{d}^{\mathrm{obs}}\mid\bm{m}\right)\propto-\frac{1}{2}\sum_{r}\|\bm{d}_{r}-\bm{d}_{r}^{\text{obs}}\|^{2}=-l_{2}(\bm{d}_{r}^{\text{obs}},\bm{d}_{r}).$ (33)
To obtain $\nabla_{\bm{m}}\log p\left(\bm{d}^{\mathrm{obs}}\mid\bm{m}\right)$, the gradient of the misfit function $l_{2}$ with respect to $\bm{m}$ is required. The gradient is given by
$\nabla_{\bm{m}}l_{2}(\bm{d}_{r}^{\text{obs}},\bm{d}_{r})=\sum_{r}\left(\bm{d}_{r}-\bm{d}_{r}^{\text{obs}}\right)\cdot\nabla_{\bm{m}}\bm{d}_{r}.$ (34)
Since $\bm{d}_{r}$ depends on the model parameters $\bm{m}$, it is necessary to compute $\nabla_{\bm{m}}\bm{d}_{r}$, which involves differentiating the forward model $\mathcal{L}$ with respect to $\bm{m}$.
To avoid direct computation of the gradient of $\bm{d}_{r}$ with respect to $\bm{m}$ (which is computationally challenging), the adjoint state method [41] is used. The adjoint state, denoted as $\bm{a}_{r}$, is an auxiliary variable introduced to propagate the error backward through the system. By solving the backward equation for the adjoint state, the gradient of the misfit function is computed without the need to differentiate the forward model
The adjoint state $\bm{a}_{r}$ satisfies a backward equation, which is derived by applying the chain rule to the misfit function $l_{2}$. The adjoint equation is
$\mathcal{L}^{\dagger}\bm{a}_{r}=\bm{d}_{r}-\bm{d}_{r}^{\text{obs}},$ (35)
where $\mathcal{L}^{\dagger}$ is the adjoint of the forward operator $\mathcal{L}$. The adjoint state $\bm{a}_{r}$ propagates the error backward from the data space to the model space. Once the adjoint state $\bm{a}_{r}$ is obtained, the gradient of the misfit function with respect to the parameters $\bm{m}$ is
$\nabla_{\bm{m}}l_{2}(\bm{d}_{r}^{\text{obs}},\bm{d}_{r})=\sum_{r}\bm{a}_{r}\cdot\nabla_{\bm{m}}\mathcal{L}(\bm{x}_{r},t,\bm{m}).$ (36)
To evaluate the gradient of $\log p(\bm{m})$, it is assumed that the prior distribution $p(\bm{m})$ follows a Gaussian distribution with mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$,
$p(\bm{m})=\frac{1}{(2\pi\det\mathbf{\Sigma})^{\frac{d}{2}}}\exp\left(-\frac{1}{2}(\bm{m}-\bm{\mu})^{\top}\mathbf{\Sigma}^{-1}(\bm{m}-\bm{\mu})\right),$ (37)
where $\mathbf{\Sigma}^{-1}$ is the inverse of the covariance matrix. To simplify the calculation, the mean-field Gaussian parameterisation is employed [33]. The prior distribution $p(\bm{m})$ is approximated by a factorized Gaussian distribution over the parameters $\bm{m}=(m_{1},m_{2},\dots,m_{d})$,
$p(\bm{m})=\prod_{i=1}^{d}p(m_{i}).$ (38)
In the above factorized form, the probability $p(m_{i})$ for each dimension is a Gaussian distribution
$p(m_{i})=\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(-\frac{(m_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (39)
where $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the $i$-th component $m_{i}$. The factorization (38) reduces the complexity of the prior distribution by assuming that the parameters are independent. The gradient of $\log p(\bm{m})$ with respect to each $m_{i}$ is
$\nabla_{m_{i}}\log p\left(\bm{m}\right)=$ $\frac{\partial}{\partial m_{i}}\left(-\frac{1}{2}\log(2\pi\sigma_{i}^{2})-\frac{(m_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right)=-\frac{m_{i}-\mu_{i}}{\sigma_{i}^{2}}.$ (40)
Stacking the gradients for all components $m_{i}$ results in the gradient $\nabla_{\bm{m}}\log p(\bm{m}\mid\bm{d}^{\mathrm{obs}})$.
4 Simulated Results
Figure 2: Inversion setup and initial conditions for the synthetic data experiment using a linear array transducer. (a) The true model setup. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The Ricker wavelet used as the source waveform. (c) The initial model for the conventional FWI. (d) An example of the initial particles for SVGD.
In this section, the gradients in Eqs. (36) and (40) are combined to apply the SVGD update rule (24) for solving FWI problems under in-silico experimental conditions. The primary focus is on the SOS distribution of objects or tissues within the acoustic field. Using the obtained SOS information, imaging of the objects or tissues is performed. To assess the convergence of the SVGD algorithm, Eq. (4) is used as the loss function. Once the update steps converge, the SOS distribution is visualized by plotting the mean of the particles, and the uncertainty in the corresponding SOS is measured through the standard deviation. Additionally, to evaluate the effectiveness of SVGD, the conventional FWI method is applied to obtain a deterministic result. In conventional FWI, the gradient is computed using the adjoint method, and the loss function (4) is minimized via gradient descent.
4.1 Linear Array Transducer
Figure 3: Results of the inversion process for the linear array transducer simulation. (a) The SOS distribution obtained from the conventional FWI. (b) The mean result of the SVGD particles. (c) The standard deviation of the SOS distribution from SVGD. (d) The loss values plotted against the iteration number.
First, the focus is on using synthetic data from a linear array transducer for inversion imaging. A two-dimensional grid model is employed, with a grid size of $201\times 201$ and a spatial step size of $0.5$ mm, resulting in a total computational area of $10\,\text{cm}\times 10\,\text{cm}$. As shown in Figure 2(a), the true model setup consists of a circular object located at the center of the computational area. The object is a homogeneous isotropic medium with a diameter of $1.5$ cm and an SOS of $1.7$ km/s. The background SOS is set at $1.5$ km/s, resembling the speed of sound under water. This difference in SOS results in variations in the wave propagation, providing a basis for the inversion algorithm to distinguish the target object from the background medium.
The simulated experiment utilizes 21 transmitting sources and 101 receivers. In Figure 2(a), the sources are evenly distributed at $x=0.15$ cm, and the receivers are uniformly placed at $x=9.9$ cm. The starting and ending positions of both sources and receivers are $0.5$ cm away from the upper and lower boundaries of the computational area, respectively. The source waveform is set as the Ricker wavelet, with a central frequency of $0.5$ MHz and a duration time of $0.2$ ms. As illustrated in Figure 2(b), the frequency spectrum of the Ricker wavelet is well-concentrated, providing high energy within a narrow frequency band.
For the conventional FWI, the initial model sets the SOS within the grid to $1.5$ km/s, as shown in Figure 2(c). For the SVGD algorithm, Gaussian white noise with a mean of 00 and a standard deviation of $0.01$ km/s is added to the initial FWI model to generate different initial particles. Figure 2(d) displays one example of the initial particles. These initial models serve as starting points, which are iteratively updated during the inversion process. The prior distribution is assumed to follow the Gaussian distribution Eq. (37), with $\mu_{i}=1.5$ km/s and $\sigma_{i}=1$ km/s. During the inversion iterations, the Devito software (https://github.com/devitocodes/devito) is used to obtain forward data and the gradient from the adjoint-state method.
Figure 3 presents the results of the inversion process. The SOS distribution obtained from the conventional FWI is shown in Figure 3(a). The circular object in the true model is reconstructed as an ellipse in the inversion result. This transformation is primarily due to insufficient angular coverage. In the case of the linear array transducer, the sources and receivers do not provide adequate angular coverage, limiting the acquisition of wave propagation data from multiple angles. As a result, the inversion algorithm is unable to accurately reconstruct the true shape of the target object. A similar transformation is observed in Figure 3(b), which shows the mean SOS distribution of the SVGD particles. Compared to the conventional FWI results, the mean SOS in SVGD particles is closer to the true model. The standard deviation, shown in Figure 3(c), measures the uncertainty in the mean SOS distribution. The central region of Figure 3(c) exhibits low uncertainty, indicating that the SVGD algorithm provides a more reliable SOS estimate in this area. Thus, the standard deviation serves to evaluate the accuracy and reliability of the mean result across different regions. Figure 3(d) plots the loss values against the iteration number, showing the convergence of the inversion process.
4.2 Ring Array Transducer
Figure 4: Results of the inversion process for the ring array transducer simulation. (a) The true model setup. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The SOS distribution obtained from the conventional FWI. (c) The mean SOS from SVGD. (d) The standard deviation of the SOS distribution from SVGD. (e) The loss values plotted against the iteration number. (f) The relative error between the predicted mean SOS and the true SOS. The gray dashed circle (radius = 3.25 cm) indicates the region used to evaluate the effectiveness of SVGD.
To alleviate the effects of insufficient angular coverage caused by the linear array transducer setup, an alternative transducer configuration commonly used in medical ultrasound imaging is the ring array. The inversion imaging process is then performed using simulated data from the ring array transducer configuration. As shown in Figure 4(a), the true model consists of a rectangular object positioned at the center of the computational domain, with both width and length measuring $3$ cm. The central region has an SOS of $1.7$ km/s, with a $1.5$ cm × $1.5$ cm rectangular core, while the surrounding region has an SOS of $1.6$ km/s. The background SOS is set at $1.5$ km/s. The experiment uses $32$ sources and $32$ receivers, with all sources uniformly distributed on a circle with a diameter of 8.5 cm, centered at the grid’s center. The layout of the receivers is identical to that of the sources. The remaining parameters are the same as those in the linear array configuration shown in Figure 2.
[…]
5 Breast Data Results
Figure 5: Inversion results for the breast tissue model with a tumor. (a) The true SOS distribution in breast tissue, incorporating acoustic tissue properties and a tumor. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b)The SOS distribution obtained from the conventional FWI. (c) The mean result of the SVGD algorithm. The magenta arrows in (a)-(c) highlight the location of the tumor. (d) The standard deviation of the SOS distribution from the SVGD algorithm.
In the previous section, the effectiveness of the SVGD algorithm is validated using both linear and ring array transducers. This section applies the method to a complex and practical medical tissue scenario. An open breast tissue database is used, incorporating acoustic tissue properties and adding a tumor, as described in [42, 43]. Figure 5(a) shows the true SOS model. The imaged object represents the SOS distribution of breast tissue, with SOS values ranging from 1.46 km/s to 1.65 km/s. The spatial discretization of the model consists of 218 x 233 grid points, with a grid spacing of 0.5 mm. The setup includes 64 sources and 64 receivers arranged in an elliptical pattern around the center of the computational domain. A Ricker wavelet with a central frequency of 0.5 MHz and a duration of 0.25 ms is used as the transmitted waveform. The remaining parameters follow the same setup as in Figure 2.
Figure 6: Loss and relative error analysis for breast tissue inversion imaging. (a) The loss values plotted against the iteration number. (b) and (c) show the relative error between the predicted and true SOS for FWI and SVGD, respectively. The gray dashed ellipse highlights the region to assess the effectiveness of the inversion algorithms.
In the inversion process, the initial model for FWI is set to a constant SOS of $1.5$ km/s. For the SVGD algorithm, different initial particles are generated by adding Gaussian white noise to the initial model. The prior distribution follows the Gaussian distribution (37) with $\mu_{i}=1.5$ km/s and $\sigma_{i}=1$ km/s. Figures 5(b) and (c) show the SOS distributions obtained by FWI and SVGD, respectively. Both methods successfully reconstruct the breast tissue information, identifying structures such as the skin, fat, stroma, and tumor. Figure 5(d) measures the imaging uncertainty of the SVGD mean SOS. The uncertainty varies across different tissues in the breast, with most values smaller than 0.023 and a mean uncertainty of approximately 0.012. The low uncertainty suggests that the SVGD mean SOS provides a reliable estimate of the breast tissue properties.
Figure 6(a) shows the loss values plotted against the iteration number for both the conventional FWI and SVGD algorithms. The FWI algorithm reaches a lower loss value after $60$ inversion steps. Figures 6(b) and (c) display the relative error between the predicted and true SOS for FWI and SVGD, respectively. Within the gray dashed ellipse, the maximum relative error for FWI is $5.714\%$, with a mean relative error of $0.514\%$. For SVGD, the maximum relative error is $5.549\%$, and the mean relative error is $0.512\%$. Despite the slightly higher loss value for SVGD, it does not imply poorer SOS estimation. Both FWI and SVGD algorithms provide accurate and reliable estimates of the breast tissue SOS distribution.
The results in Figures 5-6 validate the effectiveness of using the mean value from SVGD for reconstructing breast tissue. To further assess the quality of the obtained standard deviation, SVGD is compared with the Stochastic Variational Inference (SVI) method introduced in [33]. The SVI-based FWI assumes a prior distribution as a standard normal distribution with mean $\bm{\mu}$ and covariance $\mathbf{\Sigma}$. This method estimates the pixel-wise standard deviation of the SOS distribution using a mean-field Gaussian approximation. The gradient of the L2-norm loss with respect to $\bm{\mu}$ (or $\mathbf{\Sigma}$) is computed using the adjoint-state method. The parameter $\bm{m}$ is derived from the updated standard normal distribution. When the loss converges or the iteration reaches a predefined threshold, the mean value is used for tissue reconstruction, and the standard deviation quantifies the uncertainty associated with the imaging results.
Figure 7: SVI results for breast tissue. (a) The mean SOS from the SVI algorithm, with the magenta arrow highlighting the location of the tumor. (b) The standard deviation of the SOS distribution. (c) The loss values plotted against the iteration number. (d) The relative error between the predicted and true SOS. The gray dashed ellipse highlights the region to assess the effectiveness of the inversion algorithms.
Figure 7 (a) presents the mean SOS obtained from the SVI method, where both the breast tissue and the tumor are well reconstructed. However, the standard deviation shown in Figure 7 (b) fails to effectively quantify the uncertainty in the imaging results. In comparison to the uncertainty illustrated in Figure 5 (d), only the skin layer is clearly depicted in Figure 7 (b). The maximum standard deviation within the breast tissue is 0.14, with an average of 0.1. This standard deviation is much larger than that obtained from SVGD, suggesting that SVGD provides a tighter lower and upper bound for the prior distribution. From the variations in loss values over iterations shown in Figure 7 (c), it is evident that the SVI method achieves the converged SOS results. Figure 7 (d) shows the relative error between the predicted and true SOS for SVI. Inside the gray dashed circle, the maximum relative error is $6.121\%$, and the mean relative error is $0.558\%$. The relative error of SVI is slightly larger than that of conventional FWI and SVGD methods.
Figure 8: Variations of SOS along certain lines. (a) The two selected lines at $x=5.4$ cm and $y=5.5$ cm. (b) and (c) show variations of SOS along the lines at $x=5.4$ cm and $y=5.5$ cm, respectively.
The conventional FWI, SVGD, and SVI effectively reconstruct the breast tissue, with SVGD demonstrating the best performance among the three methods. To compare the reconstructed results in a local area, two lines are selected at $x=5.4$ cm and $y=5.5$ cm, and the differences in the inversion results are plotted. As shown in Figure 8(a), one vertical line at $x=5.4$ cm and one horizontal line at $y=5.5$ cm are used for comparison. Figure 8(b) shows the reconstructed SOS for the line at $x=5.4$ cm. The SOS obtained by SVGD (green line) is closer to the true value (red line), particularly in the regions where the SOS exceeds $1.55$ km/s, which suggests that SVGD is promising for approximating larger SOS areas in FWI imaging. Figure 8(c) presents the reconstructed SOS for the line at $y=5.5$ cm, the three methods align with the true SOS, and SVGD provides a better approximation in larger SOS regions.
Figure 9: Variations of the standard deviation from SVGD and SVI. (a) and (b) show the variations of standard deviation along the lines at $x=5.4$ cm and $y=5.5$ cm, respectively.
To analyze the uncertainty quantification obtained by the two VI-based methods, SVI and SVGD, Figures 9 (a) and (b) show the variations of the standard deviation along the two lines at $x=5.4$ cm and $y=5.5$ cm, respectively. Compared to the SVI, the variations in the standard deviation from the SVGD algorithm are more consistent with the true SOS variations, particularly in regions with larger SOS values. Additionally, the standard deviation values from SVGD are considerably smaller than those from SVI, indicating that SVGD provides a tighter lower and upper bound for the prior distribution.
6 Discussion
FWI is a powerful technique for imaging biological tissue properties, yet estimating uncertainty and solving Bayes’ equation for high-dimensional acoustic data is computationally expensive. This study tackles this challenge by using the SVGD algorithm. By combining the adjoint method with a Gaussian prior assumption, Bayes’ equation is solved to obtain the posterior distribution of the parameters. The efficiency of the SVGD algorithm in FWI depends on the choice of prior distribution. Selecting an appropriate prior can accelerate the algorithm’s ability to produce accurate results. One possible solution for acceleration is to incorporate deep learning models, such as Variational Autoencoders (VAE) [44, 45, 46] or Generative Adversarial Networks (GANs) [47, 48, 49], which learns the parameter distribution from medical imaging data like MRI or USCT. These models generate more realistic priors, further improving the performance of the SVGD algorithm in complex inversion tasks.
The SVGD algorithm offers the advantage of quantifying uncertainty in the imaged regions, allowing for better decision-making when interpreting the results. A key observation is that the lower standard deviation, the more reliable the inversion result for that region. Areas with higher standard deviation suggest that the solution may be unstable or unclear. The high uncertainty could be mitigated by giving more importance to regions with lower standard deviation. One potential strategy is to introduce an adaptive weighting scheme that assigns lower weights to noisy data and higher weights to cleaner data. This scheme is further integrated into the loss function to minimize the impact of noisy data. Besides, incorporating techniques such as stochastic gradient descent (SGD) [50, 51, 52] or mini-batch gradient estimation [53, 54, 55], the adaptive weighting scheme ensures that more important subsets of the data are used in each iteration, helping to stabilize the particle updates. In regions of high uncertainty, applying additional methods such as smoothing, noise filtering, or regularization helps refine the parameter estimates, ultimately improving the overall resolution of the FWI.
In clinical diagnosis, the standard deviation may accelerate artificial intelligence (AI) diagnosis by highlighting uncertain results caused by noise, artifacts, or insufficient data. The ability to quantify uncertainty allows AI systems to flag regions that need further examination or additional data collection. By focusing on regions with high uncertainty, AI guides clinicians in prioritizing additional imaging or confirming diagnoses. This capability helps reduce the likelihood of misdiagnosis and supports more informed clinical decision-making. By providing a reliable measure of uncertainty, the standard deviation ultimately enhances the trustworthiness of FWI-based imaging techniques, making them more effective in real-world clinical applications.
7 Conclusion
In this study, an SVGD-based FWI algorithm is employed for medical ultrasound imaging, demonstrating its effectiveness in reconstructing the SOS distributions of imaged objects. For comparison, the conventional FWI and SVI are also used to reconstruct the SOS. The results show that all three algorithms provide accurate estimates of object properties, with SVGD offering additional benefits, such as faster convergence and the ability to quantify uncertainty. The SVGD results exhibit low relative errors and minimal uncertainty in the core regions of the object. Furthermore, the performance of three algorithms is evaluated using a realistic breast tissue model. The SVGD algorithm demonstrates comparable or even superior accuracy to other two algorithms. The uncertainty quantified by SVGD aligns more closely with the true SOS variations and provides tighter bounds for the prior distribution, positioning SVGD-based FWI as a promising tool for inversion imaging tasks.
Future work focus on incorporating prior information into the SVGD algorithm. By adding covariance to the prior, the accuracy and efficiency of the reconstruction process may be further improved. Additionally, integrating shear wave data with the SVGD algorithm could enhance elasticity imaging, particularly for assessing tissue stiffness, which is critical in evaluating conditions such as musculoskeletal tissue disorders and tumors. Combining SVGD with shear wave data may also help mitigate the effects of cycle-skipping near bone tissue, a challenge often encountered in FWI. This approach may require fine-tuning and adaptation of the algorithm to address specific challenges in different scenarios, but it holds potential for enhancing the diagnostic capabilities of ultrasound and other imaging modalities.
8 Acknowledgement
This work was supported by the National Natural Science Foundation of China (grant numbers 12122403, 12034005, and 12327807). The computations for this research were performed using the CFFF platform at Fudan University. Thanks to Oscar Bates for his constructive suggestions.
9 Data availability
Data will be made available on reasonable request.